At what distance apart would two equal masses of 150 kg need to be placed for the force between them to be 2.0
x 10^-5 N?
1
Expert's answer
2020-04-15T10:42:32-0400

The Newton's gravity law says

"F=G\frac{m_1m_2}{r^2}."

Hence, the distance between masses

"r=\sqrt{\frac{Gm^2}{F}}=\sqrt{\frac{6.67\times 10^{-11}\times (150)^2}{2.0\times 10^{-5}}}=0.27\:\rm m."